A kite is flying at the end of a 300-foot string. It is 120 feet above the ground. About how far away horizontally is the person holding the string from the kite?
Answer by checkley75(3666)

120^2+x^2=300^2
14,400+x^2=90,000
x^2=90,000-14,400
x^2=75,600
x=274.95 ft. from the person holding the string.